OpenStudy (anonymous):

Factor completely: xy + 4y − 2x − 8

OpenStudy (callisto):

xy + 4y − 2x − 8 = y(x+4) -2(x+4) = (y-2)(x+4)

OpenStudy (callisto):

first take out common factor of the term consisting same variable, then, take out the common factor again